OpenStudy (anonymous):

Divide: 5x^2-20x/4x-9 divide by x-4/4x-9

OpenStudy (saifoo.khan):

\[\frac{5x^2-20x}{4x-9} \div \frac{ x-4}{4x-9 }\]Change the divide sign: \[\frac{5x^2-20x}{4x-9} \times \frac{ 4x-9}{x-4 }\]

OpenStudy (anonymous):

Ok I get that part, then what?

OpenStudy (anonymous):

factor a 5x out of the top and you get 5x(x-4)

OpenStudy (saifoo.khan):

\[\frac{5x^2-20x}{\cancel{4x-9}} \times \frac{ \cancel{4x-9}}{x-4 }\]

OpenStudy (anonymous):

Do you see something in the top and bottom that can cancel out before you do any factoring?

OpenStudy (anonymous):

@saifoo.khan just illustrated my point

OpenStudy (anonymous):

I was going to say 4x-9

OpenStudy (anonymous):

Change sign, cancel out, what is the next step

OpenStudy (saifoo.khan):